# From Specific-MLLMs to Omni-MLLMs: A Survey on MLLMs Aligned with Multi-modalities

### v2

@@ C.3 Performance of Omni-MLLMs
We statistic the performance of various mainstream Omni-MLLMs in uni-modal understanding, cross-modal understanding, and cross-modal, as shown in Table 5. We also show the performance of several Specific-MLLMs (Lin et al., 2024b; Li et al., 2024a; Chu et al., 2023; Xu et al., 2024b; Sun et al., 2024c; Jin et al., 2024a) on selected tasks for comparison. The results are mainly from corresponding papers (some results are used as baselines in other papers). It is worth noting that due to differences in the size and performance of the pre-trained models, Omni-MLLMs with the same backbone LLM may still not be fairly comparable. Therefore, this table only provides a rough trend of performance.
@@
Table 5: The performance of Omni-MLLMs on different benchmarks. The selected uni-modal understanding benchmarks include Video-Text2Text (Xu et al., 2017a, 2016), Image-Text2Text (Goyal et al., 2017; Plummer et al., 2015; Liu et al., 2024d), Audio-Text2Text (Kim et al., 2019; Lipping et al., 2022), and 3D-Text2Text (Deitke et al., 2023). The chosen uni-modal generation benchmarks include Text2Image (Lin et al., 2014), Text2Video (Xu et al., 2016), and Text2Audio (Kim et al., 2019). The selected cross-modal understanding benchmarks are Image-Audio-Text2Text (Chen et al., 2020b; AlAmri et al., 2018) and Video-Audio-Text2Text (Li et al., 2022b, a).

### v3

@@ C.3 Performance of Omni-MLLMs
We statistic the performance of various mainstream Omni-MLLMs in uni-modal understanding, cross-modal understanding, and cross-modal, as shown in Table 5. We also show the performance of several Specific-MLLMs (Lin et al., 2024; Li et al., 2024a; Chu et al., 2023; Xu et al., 2024b; Sun et al., 2024c; Jin et al., 2024) on selected tasks for comparison. The results are mainly from corresponding papers (some results are used as baselines in other papers). It is worth noting that due to differences in the size and performance of the pre-trained models, Omni-MLLMs with the same backbone LLM may still not be fairly comparable. Therefore, this table only provides a rough trend of performance.
@@
Table 5: The performance of Omni-MLLMs on different benchmarks. The selected uni-modal understanding benchmarks include Video-Text2Text (Xu et al., 2017, 2016), Image-Text2Text (Goyal et al., 2017; Plummer et al., 2015; Liu et al., 2024d), Audio-Text2Text (Kim et al., 2019; Lipping et al., 2022), and 3D-Text2Text (Deitke et al., 2023). The chosen uni-modal generation benchmarks include Text2Image (Lin et al., 2014), Text2Video (Xu et al., 2016), and Text2Audio (Kim et al., 2019). The selected cross-modal understanding benchmarks are Image-Audio-Text2Text (Chen et al., 2020b; AlAmri et al., 2018) and Video-Audio-Text2Text (Li et al., 2022b, a).